### v1

# CRAFT: Training-Free Cascaded Retrieval for Tabular QA

Abstract
Table Question Answering (TQA) involves retrieving relevant tables from a large corpus to answer natural language queries. Traditional dense retrieval models, such as DTR and ColBERT, not only incur high computational costs for large-scale retrieval tasks but also require retraining or fine-tuning on new datasets, limiting their adaptability to evolving domains and knowledge. In this work, we propose CRAFT, a cascaded retrieval approach that first uses a sparse retrieval model to filter a subset of candidate tables before applying more computationally expensive dense models and neural re-rankers. Our approach achieves better retrieval performance than state-of-the-art (SOTA) sparse, dense, and hybrid retrievers. We further enhance table representations by generating table descriptions and titles using Gemini Flash 1.5 Team et al. (2025). End-to-end TQA results using various Large Language Models (LLMs) on NQ-Tables Herzig et al. (2021), a subset of the Natural Questions Dataset, demonstrate CRAFT effectiveness.
CRAFT: Training-Free Cascaded Retrieval for Tabular QA
$**{}^{\textbf{*}}$darsh $**{}^{\textbf{*}}$ushal Raj Bhandari2  Jianxi Gao2 Soham Dan3Vivek Gupta1 1Arizona State University 2Rensselaer Polytechnic Institute 3Microsoft asing725@asu.edubhandk@rpi.edu
footnotetext: *These authors contributed equally to this work.   †These authors jointly supervised this work.
1 Introduction
Tables serve as a vital format for representing structured knowledge, yet enabling open-domain question answering (QA) systems to retrieve and interpret tabular information effectively remains a non-trivial challenge. Unlike traditional Tabular QA (TQA) settings that assume access to gold tables, open-domain QA must first identify relevant tables from a large corpus before reasoning over their structure to derive accurate answers Liang et al. (2024). Simple approaches that linearize tables into text and apply standard text-retrieval methods can lose the syntactic and structural information encoded in rows and columns Badaro et al. (2023); Min et al. (2024). As a result, effective table retrieval has been identified as a key bottleneck for scalable QA over tabular data Li et al. (2025).
Figure 1: Overview of the CRAFT Framework.
Recent work has started to tackle open-domain TQA by developing specialized retrieval models. Early approaches adopted sparse or dense text retrievers on flattened tables (e.g., BM25 or bi-encoder models)Fan et al. (2024); Robertson and Zaragoza (2009). For example, the Dense Table Retriever (DTR) introduced by Herzig et al. (2021) augments a bi-encoder with row and column features to better capture table structure. More recent methods explicitly incorporate fine-grained structure: Jin et al. (2023) propose a syntax- and structure-aware retriever that represents each question phrase and each table header and values separately, achieving state-of-the-art results on the NQ-Tables benchmark. Likewise, Li et al. (2025) introduces THYME, a field-aware hybrid retriever that tailors matching strategies to different table fields (e.g., cell values vs. captions). THYME significantly outperforms prior baselines on open-domain table retrieval, highlighting the gains from table-specific modeling. Similarly, Pan et al. (2022) introduced T-RAG, an end-to-end table QA framework that combines table retrievers with sequence-to-sequence generators for retrieval-augmented generation. These advances, however, often rely on complex architectures or extensive fine-tuning on the TQA dataset. This raises an intriguing question: Can off-the-shelf pretrained models, when integrated into a carefully designed retrieval pipeline, achieve competitive performance to expensive domain-specialized, fine-tuned methods? To this end, we propose CRAFT, a modular and scalable multi-stage retrieval framework for TQA that harnesses pretrained models without relying on dataset-specific fine-tuning.
Our work introduces CRAFT, a training free cascaded tabular retriever algorithm that achieves competitive performance with state-of-the-art tabular retrievers requiring extensive in-domain training. Through comprehensive ablation studies, we demonstrate the robustness of CRAFT across varying retrieval threshold $K$, highlighting its adaptability and stability. Building on this foundation, we develop a complete end-to-end TQA system that outperforms all existing TQA-in-the-wild , showcasing the effectiveness of our retrieval approach in real-world, open-domain settings.
2 Methodology
To create a retrieval setup without dataset-specific fine-tuning, we adopt a modular framework that leverages off-the-shelf models with multi-stage retrieval, with more complex models for each successive stage. As shown in Figure 1, CRAFT involves prepossessing followed by stage-wise table retrieval using increasingly expressive models, culminating in answer generation via a pretrained LLM. Our approach demonstrates that pretrained models can effectively retrieve relevant tabular data and generate accurate answers compared to the state-of-the-art framework Zhao et al. (2024). The pipeline consists of leveraging preprocessing followed by cascaded table retrieval and end-to-end answer generation to maximize semantic alignment between the input queries and tabular content using off-the-shelf embeddings models.
2.1 Preprocessing
The preprocessing step involved three key operations to enhance the query and table representations for effective retrieval. Firstly, Gemini-1.5-Flash Team et al. (2025) was employed to generate sub-questions derived from each original query, thus decomposing complex queries into more granular components. Secondly, table titles and comprehensive descriptions were generated using Gemini-1.5-Flash, enriching contextual understanding and enabling more precise downstream retrieval. Finally, all rows across every table were ranked using a Sentence Transformer (all-mpnet-base-v2) model SentenceTransformers (2022), applied directly to the original queries to assess semantic relevance. The NQ-Tables Dataset statistics are shown in Table 7.
2.2 Table Retrieval
The table retrieval step involves a rigorous, three-stage filtering methodology to identify highly relevant tables for each query. In the first stage, SPLADE Formal et al. (2021) was applied to a comprehensive data set consisting of 169,898 tables, leveraging generated table titles, column headers, individual cell values, and detailed table descriptions to filter and narrow down the candidate tables to the top 5,000 entries. SPLADE leverages sparse lexical expansion to efficiently score candidate passages using minimal memory, making it well-suited for handling large-scale retrieval tasks.
In the second stage, the Sentence Transformer (all-mpnet-base-v2) provides a fine-grained semantic refinement by representing each table with mini-tables created through appending column headers to each of the top 5 rows. This refinement approach allows the selection of the top 1,000 mini-tables. In the final stage, the text-embedding-3-small OpenAI (2024) model performs a detailed re-ranking on the top 100 mini-tables, carefully assessing these mini-tables based on their semantic embeddings. Concurrently, we enrich all 919 unique queries using sub-question decomposition generated by Gemini-1.5-Flash to maximize their semantic alignment with the selected tables.
2.3 End-to-End Answer Generation
For the end-to-end table question-answering, we use the final retrieval result on the set of 919 queries with three different pretrained models: Llama3-8B, Qwen2.5-7B, and Mistralv0.2-7B. Each LLM generates answers from prompts that include a varying number of tables ($n=\{1,3,5,8,10\}$). Each prompt contains the table title, column headers, and mini-tables (top 5 rows) for $n$ tables, offering comprehensive yet concise context to support accurate answer generation. In addition, we also experiment with the inclusion of a few-shot prompt to guide the text generation process.
We evaluate model performance and accuracy by computing the F1 score, which quantifies the overlap between generated answers and reference answers, enabling comparative analysis across different models and table settings.
Model R@1 R@10 R@50
Sparse Retrieval Models
BM25 18.49 36.94 52.61
SPLADE 39.84 83.33 94.65
Dense Retrieval Models
BIBERT 43.78 82.25 93.71
DPR 45.32 85.84 95.44
TAPAS 43.79 83.49 95.10
DTR 32.62 75.86 89.77
SSDR(im) 45.47 84.00 95.05
T-RAG* 46.07 85.40 95.03
Hybrid Retrieval Models
BIBERT+BM25 35.87 79.63 94.56
DHR 43.67 84.65 95.62
BIBERT+SPLADE 45.62 86.72 95.62
THYME 48.55 86.38 96.08
CRAFT Stage-wise
Stage 1 34.38 72.90 91.62
     + Stage 2 36.65 82.91 96.08
        + Stage 3 41.13 87.16 96.84
CRAFT 41.13 87.16 96.84
Table 1: Table Retrieval Performance on NQ-Tables. R denotes Recall Score. Best performances are in Bold. All models except CRAFT are trained on NQ-Tables. ‘*’ denotes results reported for the full test-set with all 966 queries.
3 Results and Analysis
CRAFT Retrieval Performance.
Table 1 presents the retrieval performance of three groups of methods: Sparse, Dense, and Hybrid Retrievers on the NQ-Tables dataset. Hybrid retrievers show the best overall performance, and dense retrievers outperform Sparse Retrievers. Our method achieves performance comparable to SOTA fine-tuned retrievers, even outperforming them at R@10 and R@50. THYME achieves the best average performance among all retrievers. Compared to SPLADE and DTR our three-stage cascade improves retrieval step by step. In Stage 1, SPLADE finds almost all relevant tables at higher ranks (R@50 = 91.62%) but only 34.38% at the top rank (R@1). In Stage 2, the Sentence-Transformer reranker boosts mid-level performance, increasing R@10 by over 10 percentage points and pushing R@50 above 96%, while still keeping strong overall recall. Finally, the neural reranker in Stage 3 focuses on improving top results, raising R@1 by another 4.5 points and refining R@10 to 87.16%. Although CRAFT achieved a lower recall score at R@1, we observed that for some queries, there are tables in the corpus, other than the gold table, that also contains the answer to the query.
Robustness of CRAFT Retrieval.
Model R@1 R@10 R@50 $\Delta$ (Avg.)
Original Query
DTR (M) 38.74 75.73 88.36 -
DTR (M) + HN 47.33 80.96 91.51 -
DTR (S) 39.28 73.88 87.60 -
DTR (S) + HN 43.53 77.80 90.21 -
CRAFT 41.13 87.16 96.84 -
Perturbed Query
DTR (M) 27.09 66.27 84.33 - 8.38
DTR (M) + HN 37.87 76.06 88.47 - 5.80
DTR (S) 25.25 61.81 78.24 - 11.82
DTR (S) + HN 35.04 70.73 85.85 - 6.64
CRAFT Stage-wise
Stage 1 33.40 68.22 87.05 - 3.41
+ Stage 2 34.05 73.18 88.92 - 6.5
        + Stage 3 41.02 86.83 96.08 - 0.04
Table 2: Performance of DTR vs CRAFT on perturbed queries. M and S denote Medium and Small model sizes respectively. $\Delta$ is the Avg. change in Recall relative to the original query scores, positive values indicate improvement, and negative values indicate degradation. HN represents a model trained with Hard Negatives.
Retriever n=1 n=3 n=5
Mistral Llama3 Qwen Mistral Llama3 Qwen2.5 Mistral Llama3 Qwen2.5
BIBERT 32.93 32.66 34.80 34.6 33.16 37.92 35.30 34.28 37.01
SPLADE 29.61 32.07 31.90 35.42 37.17 35.79 34.95 33.88 37.35
SSDR_in 32.76 34.25 33.42 36.95 38.14 37.14 36.60 37.89 33.92
BIBERT-SPLADE 32.67 32.66 33.24 33.59 33.66 36.76 35.71 33.92 37.02
THYME 35.48 36.14 37.28 37.59 39.16 40.28 39.29 39.41 41.20
CRAFT 39.65 39.29 33.33 46.04 44.06 37.70 45.09 45.98 41.47
CRAFT (5-fewshot) 41.17 41.72 40.36 48.50 39.39 43.65 47.70 39.27 45.81
Table 3: Performance comparison (F1) across different retrievers, LLMs, and values of $n$ with NQ-Tables. The evaluated models include Mistral-7B, Llama3-8B, and Qwen2.5-7B, corresponding to Mistral, Llama3, and Qwen, respectively. Bolded values indicate top results.
To assess robustness under query paraphrasing, we generated perturbed versions of each question using Gemini 2.5 Flash Team et al. (2025), while keeping the ground-truth answer unchanged. Table 2 reports recall scores for DTR, trained on NQ-Tables with and without hard negatives, and our three-stage CRAFT pipeline. We directly use publicly released fine-tuned DTR models for this experiment. With the original queries, CRAFT already outperforms the DTR baselines. When using the perturbed queries, both DTR Medium and Small suffer substantial average drops of –8.38 and –11.82 points respectively. The DPR version trained with hard negatives reduces but does not eliminate this degradation (avg = –5.80 for DTR (M) and –6.64 for DTR (S)). By contrast, CRAFT’s full three-stage pipeline nearly fully recovers its original performance with a slight decrease under paraphrasing.
Ablations Analysis.
Ablation over stages further shows incremental gains: Stage 1 drops performance under perturbation by 3.4 %, Stage 2 partly recovers it, and Stage 3 restores the original recall, as shown in Appendix A.2. These results highlight CRAFT’s superior robustness to query rephrasing compared to DTR.
End-to-End Results.
We evaluate the end-to-end performance of CRAFT across multiple pretrained language models and retrieval configurations. As shown in Table 3, CRAFT consistently outperforms several state-of-the-art baselines across different models: Mistral Jiang et al. (2023), LLaMA3 Grattafiori et al. (2024), and Qwen 2.5 Qwen et al. (2025), for varying numbers of retrieved tables ($n={1,3,5}$). These results highlight the robustness of our modular, training-free approach.
When considering only one retrieved table ($n=1$), CRAFT notably outperforms baseline methods, significantly improving accuracy scores using all the pretrained models. Moreover, incorporating a 5-shot prompt further enhances performance substantially across all models and top-k tables. These results highlight that structured, cascaded retrieval by itself outperforms systems that rely on dataset‑specific training, and that augmenting this retrieval with lightweight few‑shot prompting delivers consistent additional gains. We further explore the impact of performance using different few-shot prompting and the effect on context reduction with CRAFT in Section B.
Model n =1 n=5
RAG Lewis et al. (2020) - 39.67
DPR-RAGE Lin et al. (2023) - 49.68
LI-RAGE Lin et al. (2023) - 54.17
TABERT
   - Mistral-Small-Instruct 50.26 57.14
   - Llama-3.1-70B-Instruct 49.75 56.94
   - Qwen2.5-72B-Instruct 48.75 49.17
   - Llama-3.3-70B-Instruct 46.21 55.16
Table 4: F1 scores of CRAFT with larger LLMs compared to dataset-specific baselines for $n=1$ and $n=5$ retrieved tables. Bold indicates the best performance.
Table 4 further emphasizes the effectiveness of CRAFT, particularly when employing larger, instruction-tuned LLMs. Compared to prior dataset-specific baselines, including RAG Lewis et al. (2020), DPR-RAGE, and LI-RAGE Lin et al. (2023). CRAFT consistently delivers superior F1 scores without requiring any dataset-specific training or finetuning. We find that using just a single retrieved table ($n=1$), Mistral-Small-Instruct achieves an F1 score of $50.26$, already surpassing the best-performing baseline, LI-RAGE ($54.17$), even though LI-RAGE benefits from deeper retrieval ($n=5$). With increased retrieval depth ($n=5$), Mistral-Small-Instruct further improves to $57.14$, highlighting significant gains from additional context. Similarly, Llama-3.1-70B-Instruct achieves a notable score of $56.94$ at $n=5$, clearly outperforming current SOTA. These results underscore that CRAFT’s structured, training-free retrieval methodology not only outperforms established baselines but also scales effectively with increased retrieval depth and few-shot prompting. This highlights the strength of CRAFT’s modular design in enabling adaptable, efficient, and high-performing TQA without needing any dataset-specific fine-tuning.
4 Conclusion
We present CRAFT, a modular, training-free, multi-stage cascaded retrieval framework for TQA that matches or exceeds the performance of state-of-the-art systems trained on NQ-Tables. Without any dataset-specific training, it achieves competitive SOTA recall scores, maintains robustness under query perturbation, and delivers a high end-to-end performance score when paired with existing pretrained LLMs. By decoupling retrieval from domain-specific fine-tuning, CRAFT offers a scalable, adaptable solution for dynamic tabular domains where rapid deployment and data scarcity are critical issues.
5 Limitations
While our evaluation focuses on the NQ-Tables benchmark, future validation on additional TQA datasets would help further assess the generalizability of our approach. Although our multi-stage retrieval pipeline, combining sparse filtering, dense semantic reranking, and neural reranking, introduces added inference steps, this design enables significantly improved retrieval accuracy, offering a trade-off between performance and runtime that can be optimized for specific deployment needs.
Appendix A CRAFT: Retrieval Analysis
A.1 Stage-wise Results
[…]
Model R@1 R@10 R@50 R@100 R@1000 R@5000
Stage 1 34.38 72.90 91.62 94.89 98.91 99.56
     + Stage 2 36.65 82.91 96.08 97.82 98.91 -
        + Stage 3 41.13 87.16 96.84 - - -
CRAFT 41.13 87.16 96.84 - - -
Table 5: Table Retrieval Performance on NQ-Tables with various stages of CRAFT. R denotes Recall Score. Best performances are in Bold. ‘*’ denotes results reported for the full test-set with all 966 queries.
A.2 Ablation Study
Stage 1: Sparse Retrieval with Context Enrichment
We test the effect of table descriptions and query decomposition using BM25 and SPLADE. The results for Stage 1 are shown in Table 6 and Table 8. Both retrievers exhibit improvement when employing Question Decomposition. Specifically, SPLADE achieves gains of 2.18% and 4.03% at R@10 and R@50 respectively. However, incorporating Table Descriptions does not yield additional improvements for SPLADE, likely because the tables alone already exhaust the maximum context length of the SPLADE model, rendering extra descriptions redundant. Notably, SPLADE achieves a Recall@5000 of 99.59%, indicating performance saturation and leaving minimal scope for further enhancement. In contrast, BM25 demonstrates substantial improvement with the inclusion of table descriptions, clearly benefiting from context-enriched lexical matching.
Stage:2 Dense Semantic Reranking and Irrelevant Context Reduction
We apply all-mpnet-base-v2 to rerank the top 5000 tables using dense embeddings. Since dense embedding models have limited context length, we rank the rows of each table and retain the 5 most relevant rows per table. Table 9 show recall gains from filtering irrelevant rows, with R@10 rising from 79.21 to 82.91 and R@50 from 94.66 to 96.08. Dense scoring better captures semantic relevance missed in Stage 1.
Stage:3 Neural Reranking
Finally, we use text-embedding-3-small to rerank the top 300 and 100 tables. Table 10 shows the reranking results, Even without fine-tuning, neural reranking significantly sharpens final table rankings.
Table R@1 R@10 R@50 R@1000 R@5000
BM25
w/o Desc 17.95 34.38 47.44 70.73 82.48
+ Desc 22.74 45.16 62.57 86.61 93.14
SPLADE
w/o Desc 32.64 70.62 87.59 98.91 99.67
+ Desc 32.64 70.62 87.59 98.91 99.67
Table 6: Stage 1: Effect of Table Descriptions on Retrieval Performance over NQ-Tables. R denotes Recall Score. N1=169,898 Tables in Corpus. Best performances are in Bold.
Train Test
Total Query Count 9,594 966
Unique Query Count 9,594 919
Avg. # Words 8.94 8.90
Table Count 169,898 169,898
Avg. # Rows 10.70 10.70
Avg. # Columns 6.10 6.10
# Golden Tables / Query 1.00 1.05
Table 7: Statistics of the NQ-Tables benchmark. Test set contains 966 queries, out of which only 919 are unique.
Question R@1 R@10 R@50 R@1000 R@5000
BM25
Original 22.74 45.16 62.57 86.61 93.14
+ SubQues 24.15 53.54 71.38 91.08 96.84
SPLADE
Original 32.64 70.62 87.59 98.91 99.67
+ SubQues 34.38 72.90 91.62 98.91 99.56
Table 8: Stage 1: Effect of Query Decomposition- Retrieval Performance on NQ-Tables. R denotes Recall Score. N1=169,898 Tables in Corpus. Best performances are in Bold.
Table Text R@1 R@10 R@50 R@100 R@1000
Original* 37.64 79.76 94.45 97.82 99.45
     – Description 34.82 79.21 94.66 97.39 99.45
        – Irrelevant Rows 36.65 82.91 96.08 97.82 98.91
Table 9: Stage 2: Effect of Irrelevant Context Reduction- Retrieval using all-mpnet-base-v2 on NQ-Tables. Original* is title, header, all cell values, and description. N2=5000 Tables in Corpus. Best performances are in Bold.
N3 R@1 R@10 R@50
300 39.03 86.02 96.89
100 41.13 87.16 96.84
Table 10: Stage 3: Effect of Neural Reranking: Retrieval Performance using text-embedding-3-small model on NQ Tables. N3 is number of tables in corpus. Best performances are in Bold.
Configuration Parameter Value
Batch Size 144
Learning Rate $1\times 10^{-5}$
Training Epochs 50
Optimizer Adam
Initialization BERT-base
THYME Dropout ($p_{sem},p_{lex}$) 0.15
THYME FLOPS Reg. ($\lambda_{q},\lambda_{t}$) $1\times 10^{-4}$
Table 11: Training and inference configuration used for all baselines.
Appendix B Further End-to-End Analysis
Retriever n=1 n=3 n=5 n=8 n=10
Mistral Llama3 Qwen Mistral Llama3 Qwen2.5 Mistral Llama3 Qwen2.5 Mistral Llama3 Qwen2.5 Mistral Llama3 Qwen2.5
no_fewshot 39.65 39.29 33.33 46.04 44.06 37.70 45.09 45.98 41.47 43.78 44.49 49.99 42.58 44.12 50.69
fewshot_1 40.43 37.54 37.94 48.28 37.35 44.69 46.11 38.27 45.07 44.85 40.40 45.46 42.39 39.94 45.05
fewshot_2 41.56 42.73 38.98 49.29 37.77 43.16 46.69 36.83 46.02 45.35 40.54 45.03 43.53 39.35 44.86
fewshot_3 40.70 39.41 40.68 48.60 36.90 44.27 47.43 36.94 45.59 44.83 39.97 44.89 43.08 38.87 44.77
fewshot_4 41.58 40.96 40.99 48.77 37.41 43.86 47.59 36.72 46.19 45.56 40.55 44.64 43.72 40.06 43.37
fewshot_5 41.17 41.72 40.36 48.50 39.39 43.65 47.70 39.27 45.81 44.94 41.27 44.55 42.31 41.07 44.21
Table 12: F1 performance scores of different LLMs (Mistral-7B, Llama3-8B, and Qwen2.5-7B) across various retrieval configurations using CRAFT. Results are shown for increasing numbers of retrieved tables ($n=1,3,5,8,10$) and different few-shot prompting strategies. Each row represents a different few-shot variant, with no_fewshot indicating zero-shot performance.
End-to-End Performance Across Multiple Fewshot Examples.
To evaluate the scalability and robustness of CRAFT, we measure F1 performance of three pretrained LLMs, Mistral-7B, Llama3-8B, and Qwen2.5-7B, under zero-shot and few-shot prompting as the number of retrieved tables increases ($n=1,3,5,8,10$). As shown in Table 12, all models exhibit a clear upward trend in F1 score with deeper retrieval, demonstrating that structured multi-stage retrieval alone substantially narrows the gap with dataset-specific fine-tuning. However, incorporating a small number of in-context examples does not consistently yield significant improvements across models. Moreover, performance gains taper beyond $n=5$, suggesting an optimal retrieval depth that balances information richness with the risk of noise accumulation. This effect becomes more pronounced as both $n$ and the number of few-shot examples increase, leading to greater performance variability.
Model n Sub-Table Full Table
Llama-3.1-8B-Instruct 1 362.70 1782.80
3 958.79 4841.75
5 1561.34 8231.99
8 2544.36 12831.89
10 3213.31 15798.45
Mistral-7B-Instruct-v0.2 1 476.58 3314.03
3 1305.60 10183.81
5 2158.50 18252.36
8 3566.04 29321.20
10 4537.62 36191.87
Qwen2.5-7B-Instruct 1 393.68 1951.08
3 1060.45 5319.05
5 1731.25 9009.43
8 2824.83 14046.00
10 3568.37 17292.88
Table 13: Comparison of average token counts between sub-table and full-table contexts across models and top-$k$ values.
Token Consumption in Sub-Table vs. Full-Table Contexts.
While deeper retrieval improves accuracy, computational efficiency remains critical for practical deployment. Table 13 compares the average token counts for each model when using only the top-$k$ sub-table context versus concatenating the entire table. Across all three models and $k\in{1,3,5}$, sub-table inputs reduce token length by over 70%. For example, Llama3 drops from 1782.80 tokens at $k=1$ to 362.70; Mistral from 3314.03 to 476.58; and Qwen2.5 from 1951.08 to 393.68. This substantial reduction directly lowers inference costs and improves runtime efficiency. By selectively presenting only the most row table fragments, CRAFT retains essential semantic content while avoiding the diminishing returns and latency overhead associated with full-table encoding.
B.1 Prompt Used For Experiments
The prompt used for end-to-end question answering in Figure 2, the table description prompt in Figure 3 and the question decomposition prompt is shown in Figure 4.
Figure 2: Prompt used for End-to-End question answering.
Figure 3: Prompt used for Table Decomposition.
Figure 4: Prompt used for Query Expansion.